# From Loops to Oops: Fallback Behaviors of Language Models Under Uncertainty

### v1

@@ Abstract
Large language models (LLMs) often exhibit undesirable behaviors, such as hallucinations and sequence repetitions. We propose to view these behaviors as fallbacks that models exhibit under uncertainty, and investigate the connection between them. We categorize fallback behaviors — sequence repetitions, degenerate text, and hallucinations — and extensively analyze them in models from the same family that differ by the amount of pretraining tokens, parameter count, or the inclusion of instruction-following training. Our experiments reveal a clear and consistent ordering of fallback behaviors, across all these axes: the more advanced an LLM is (i.e., trained on more tokens, has more parameters, or instruction-tuned), its fallback behavior shifts from sequence repetitions, to degenerate text, and then to hallucinations. Moreover, the same ordering is observed throughout a single generation, even for the best-performing models; as uncertainty increases, models shift from generating hallucinations to producing degenerate text and then sequence repetitions. Lastly, we demonstrate that while common decoding techniques, such as random sampling, might alleviate some unwanted behaviors like sequence repetitions, they increase harder-to-detect hallucinations.
From Loops to Oops: Fallback Behaviors of Language Models
Under Uncertainty
Maor Ivgi  Ori Yoran  Jonathan Berant  Mor Geva Blavatnik School of Computer Science, Tel Aviv University
@@ 1 Introduction
Figure 2: Larger Pythia models resort to more complex fallback behaviors. Models with larger parameters count produce more correct facts (green) and hallucinations (orange) while less repeating facts (blue). The green line indicates the number of ground truth answers.
In this work, we propose that the behaviors illustrated in Figure 1 can be viewed collectively as fallback behaviors, which emerge when the model faces uncertainty. We aim to categorize and analyze the relationship between these behaviors across a range of LLMs. To this end, we create controlled yet natural settings to manage the level of uncertainty, force the generation of factual information, and evaluate the correctness of each prediction. We then test three model families—Pythia (Biderman et al., 2023), Llama 2 (Touvron et al., 2023), Llama 3 (AI, 2024), and OLMo (Groeneveld et al., 2024)—considering various factors that could influence the emergence of different fallbacks: (a) number of parameters, (b) number of pretraining tokens, (c) instruction-following training, and (d) decoding algorithms. We observe a clear ordering in the appearances of different fallbacks, as demonstrated in Figure 2, which persists across all the above factors. Our findings points to repetitions as the simplest fallback, followed by degenerate text, and finally hallucinations as the most complex behavior.
We present evidence that increasing the strength of a model tends to shift its fallback behavior to more complex forms, whereas weaker models rely on simpler behaviors like repetitive text. We demonstrate that the so-called strength of a model can be the result of increasing the model’s parameter count, additional pretraining, or the addition of an instruction-following training phase. Moreover, we show that as the generation length grows and the model struggles to generate additional factually correct responses or convincing hallucinations, it shifts backwards in the fallback behaviors spectrum from hallucinations to repetitions.
We further examine the effect of the decoding scheme, and demonstrate that while random sampling may alleviate degenerate text (Holtzman et al., 2020), it increases the rate of hallucinations, which are harder to detect and potentially more damaging to the user. Despite evidence suggesting that models may be partially aware of their knowledge gaps (Kadavath et al., 2022), models are not easily steered away from non-factual generations. Even when provided with demonstrations on how to abstain rather than hallucinate, models continue to produce hallucinations.
Overall, our study shows that while increasing models’ strength improves their accuracy (Kaplan et al., 2020), the root causes of fallback behaviors persist even in the largest and best-trained models. Moreover, superficial patches such as decoding schemes fail to solve the problem and can trade easily-detected degenerate text for more nuanced and potentially harmful hallucinations.
@@ 2 Fallback Behaviors
Since their emergence as powerful generative tools (Radford et al., 2019; Brown et al., 2020), there have been numerous reports of unwanted behaviors of LLMs. Holtzman et al. (2020) showed that greedy decoding methods can cause degenerate generations, such as incoherent text or repetitions. They propose the use of nucleus sampling to mitigate this problem. In another study, Nasr et al. (2023b) demonstrate that when given out-of-distribution inputs, models might output their training data verbatim. More recently, growing evidence (Zhu et al., 2024; Zhang et al., 2023; Aichberger et al., 2024, inter alia) suggests that even the best models often generate convincing yet factually incorrect text, typically referred to as “hallucinations” or “confabulations” (Ji et al., 2022; Bang et al., 2023). Xiao and Wang (2021) investigate model uncertainty, defining it as the probabilistic uncertainty in predicting the next token, which is influenced by the model’s training and knowledge, and identify it as a cause of hallucinations. [^2] While the model may experience internal uncertainty when predicting the next token, this might not manifest in the final logits. Once a fallback behavior is selected, it may promote the token, resulting in very high logit value.
In this paper, we propose to view these seemingly independent phenomena as fallback behaviors that models adopt when uncertain, and we investigate the relationships between them. Furthermore, we hypothesize that the strength of an LLM influences its fallback behavior, and aim to understand what factors determine how a model would behave in cases of uncertainty.
Specifically, we focus on the following phenomena: 1. Sequence repetition When models face an inability to produce plausible continuations, they tend to repeat previously generated sequences, which are known to be plausible within the current context. 2. Degenerate text As shown by Holtzman et al. (2020), models can generate degenerate text that is not strictly repetitive but follows a consistent pattern, such as enumerating or repeating sentences with variations in subject entities or attributes. 3. Hallucinations In cases that require the recollection of facts which the model cannot produce, it may fabricate coherent and seemingly plausible yet factually incorrect or unfounded content. We note that creating convincingly plausible incorrect facts is not an easy feat; even in humans, this ability develops at a late stage of development (Talwar et al., 2007; Evans and Lee, 2013).
In our analysis, we additionally looked for cases of verbatim recollection of training samples Nasr et al. (2023a), but found no evidence of such fallback behavior in our setting (see further discussion in Section 6.1).
@@ 3 Experimental Setting
To directly investigate the relation between fallback behaviors and their contributing factors, our setup exposes models to naturally-occurring controlled cases of uncertainty while recalling facts.
@@
Table 1: Example questions and answers for our main datasets.
We use the following datasets:
  1. 1.
    TriviaFacts (Triv): We manually curate 95 high-quality domain-diverse questions with a list of answers (list questions) that (a) contains multiple elements, but no more than 15, (b) requires only knowledge that appears very frequently in the web, and (c) includes short elements without multiple correct phrasings. Due to its high-quality and highly controllable setup, we base the bulk of our experiments on this dataset, with minor modifications described inline as relevant to each experiment.
  2. 2.
  3. 3.
    Qampari (Qamp): We sample 100 questions from the dataset introduced by Amouyal et al. (2023), which contains naturally occurring list questions with answers from Wikipedia. [^3] While similar to our TriviaFacts dataset, this dataset is known to have non-exhaustive answer sets and different possible phrasings for the same answers, which can result in correct answers being flagged as hallucinations.
  4. 4.
    FakeQampari (FQamp): We replace the subjects from Qampari with made-up names, verifying they do not exist, forcing uncertainty.
Table 1 provides example questions. For additional details on the datasets, see Appendix A.
Generating predictions
Evaluation metrics
Given a generated output, we first parse it into a set of facts, and evaluate each one independently as correct, repeated or hallucinated. For the list questions datasets (Triv, Qamp and FQamp), this parsing is mostly straightforward as the generations are structured as lists and the ground truth is given as a set of answers. For open-ended generation, we use FactScore Min et al. (2023a) to extract atomic facts and verify them against the entities’ Wikipedia entries. As models frequently continue generating tokens after the completion of the instruction prompt, we further detect what the model did at that point. Namely, we consider the following options: 1) generating EOS token, 2) changing the topic, for example, by creating a new list/biography which we refer to as topic change or 3) continuing to predict tokens indefinitely (until the token budget is exhausted) within the same sentence/paragraph of the answer which we note by bad format. For additional information on the parsing process and example generations and endings, see Appendix B.
Models
We perform our experiments on a variety of model families, sizes, pretraining corpora sizes and finetuning stages. We evaluate both the base and chat-specific checkpoints of Llama 2 which were finetuned on instruction and dialogue data, as well as on the recently introduced Llama 3 models Touvron et al. (2023); AI (2024). Secondly, we use OLMo Groeneveld et al. (2024), which comes in multiple model sizes, includes intermediate checkpoints throughout the pretraining phase and also offers instruction-tuned variants. Finally, we make use of the Pythia model family Biderman et al. (2023) which comes in 8 different scales, each with 154 intermediate pretraining checkpoints. We also include Dolly models which are instruction-tuned Pythia checkpoints (Conover et al., 2023). This suite of models allows us to control for different factors, such as pretraining data, model scale and training procedure, examining the effect of each factor on the emergence of fallback behaviors.
4 Fallback Behaviors of Different LLMs
Takeaway: Increasing model strength through additional pretraining, more parameters, or instruction-tuning shifts from simple fallback behaviors to more complex ones, i.e., from repetitions to hallucinations.
Scaling up models and training data improves performance and reduces undesired artifacts (Kaplan et al., 2020; Brown et al., 2020). Incorporating instruction-following phases aligns LLMs’ outputs with human preferences (Ouyang et al., 2022). In this section, we investigate how these improvements influence model behavior under uncertainty. We first focus on the trade-off between sequence repetitions and hallucinations, measured as discrete shifts in fallback behaviors. Due to the various forms and broad definition of degenerate text, measuring its appearance is not straightforward. We revisit this in depth in Section 6, demonstrating that the fallback shift is a continuous rather than discrete process.
@@ 4.1 Model scale dictates the fallback behavior
Figure 3: Larger models hallucinate instead of abstaining. When completing a list of facts about fictitious entities, larger Pythia models produce more hallucinations (orange), while smaller models repeat facts (blue). Models never abstain from providing incorrect facts.
@@ 4.2 Models shift fallbacks during pretraining
Figure 4: Models that train longer shift to complex fallbacks. The more training tokens Pythia-6.9B checkpoints see (in billions), the more hallucinations they produce and fewer repetitions they generate. The green line marks the number of ground truth answers.
@@ 4.3 Instruction finetuning shifts behaviors
Repeating the experiments from Section 4.1 and comparing models to their instruction-tuned counterparts, we see a similar shift in fallback behavior; instruction-tuned models generate fewer repeated sequences and more hallucinations on average (Figures 21 and 22). For the OLMo and Llama 2 family which had undergone more finetuning than Dolly checkpoints, the results are much more pronounced, with the hallucinations portion almost doubling between scales. While pretrained LLMs are generally unable to generate sequences shorter than what they encountered during training, instruction-tuned models are more likely to produce an EOS token, preempting the generation early and resulting in fewer facts. However, we also note that while instruction finetuning can improve alignment with human preferences, it can sometimes cause finetuned models to diverge more frequently, resulting in some bad format generations.
@@ 4.4 Similar trends in open-ended generation
Figure 5: Larger Pythia models hallucinate more when generating biographies. When producing biographies of rare entities, larger models generate more atomic facts, with an increasing rate of hallucinations.
To mimic real-world user requests, we use the BioGeneration dataset, sampling completions from a subset of the Pythia model scales with a temperature of $\tau=0.5$. We use FactScore (Min et al., 2023b) to parse each generated biography into atomic facts and let ChatGPT 3.5 (Ouyang et al., 2022) verify them against Wikipedia entries. [^4] While FactScore, which relies on ChatGPT 3.5, is not fully robust and can miss some atomic facts or incorrectly label them, we assume such errors occur at similar rates across generations, resulting in reliable trend analysis. Without the limitation on number of facts to produce, we observe that the larger models also generate more facts. Interestingly, when averaging over the very-rare entities (Figure 5) we see that even in this natural scenario, when models are required to elaborate on topics they know little about, they fall back to the same behaviors, with shifts occurring as predictably as in the controlled settings. Similar trends emerge over the rest of the popularity levels, though the more frequent an entity is, the more likely the models are to be able to recall facts for them and less uncertainty they face, thus making the results less useful for our analysis (Figures 27, 28, 29, 30 and 31). This aligns only partially with Min et al. (2023b), who found stronger models generate more atomic facts and struggle more with rare entities. However, they observed that within a model family, larger models are generally more precise (i.e., hallucinate less). In contrast, Lin et al. (2022) found the largest models are the least truthful, which aligns with our findings.

### v2

@@ Abstract
Large language models (LLMs) often exhibit undesirable behaviors, such as hallucinations and sequence repetitions. We propose to view these behaviors as fallbacks that models exhibit under epistemic uncertainty, and investigate the connection between them. We categorize fallback behaviors — sequence repetitions, degenerate text, and hallucinations — and extensively analyze them in models from the same family that differ by the amount of pretraining tokens, parameter count, or the inclusion of instruction-following training. Our experiments reveal a clear and consistent ordering of fallback behaviors, across all these axes: the more advanced an LLM is (i.e., trained on more tokens, has more parameters, or instruction-tuned), its fallback behavior shifts from sequence repetitions, to degenerate text, and then to hallucinations. Moreover, the same ordering is observed during the generation of a single sequence, even for the best-performing models; as uncertainty increases, models shift from generating hallucinations to producing degenerate text and finally sequence repetitions. Lastly, we demonstrate that while common decoding techniques, such as random sampling, alleviate unwanted behaviors like sequence repetitions, they increase harder-to-detect hallucinations.
@@ 1 Introduction
In this work, we propose that the undesired behaviors illustrated in Figure 1 can be viewed collectively as fallback behaviors, which emerge when the model faces epistemic uncertainty, namely uncertainty due to lack of knowledge (Hou et al., 2023). We aim to categorize and analyze the relationship between these behaviors across a range of LLMs. To this end, we create controlled yet natural settings, which force the generation of factual information while introducing varying levels of model uncertainty by construction. We then test three model families—Pythia (Biderman et al., 2023), Llama 2 (Touvron et al., 2023), Llama 3 (Meta AI, 2024), and OLMo (Groeneveld et al., 2024)—considering various factors that could influence the emergence of different fallbacks: (a) number of parameters, (b) number of pretraining tokens, (c) instruction-following training, and (d) decoding algorithms. We observe a clear ordering in the appearances of different fallbacks, as demonstrated in Figure 2 and persisting across all the above factors: repetitions are the simplest fallback, followed by degenerate text, and finally hallucinations as the most complex behavior.
We present evidence that increasing the strength of a model [^2] We consider prolonged training time, increased parameter count and inclusion of instruction tuned phase as “increasing models’ strength” as these techniques are ubiquitous in creation of more advanced LLMs. tends to shift its fallback behavior to more complex forms, whereas weaker models rely on simpler behaviors like repetitive text. We demonstrate that the so-called strength of a model can be the result of increasing the model’s parameter count, additional pretraining, or the addition of an instruction-following training phase.
@@
Moreover, we show that as generation length grows and the model struggles to generate factually correct responses or convincing hallucinations, it shifts back in the fallback behaviors spectrum from hallucinations to repetitions.
We further examine the effect of the decoding scheme, demonstrating that while random sampling may alleviate degenerate text (Holtzman et al., 2020), it increases the rate of hallucinations, which are harder to detect and potentially more damaging to the user. Despite evidence suggesting models may be partially aware of their knowledge gaps (Kadavath et al., 2022), they are not easily steered away from non-factual generations. Even when shown how to abstain rather than hallucinate, models continue to produce hallucinations.
Our study shows that although increasing model size reduces epistemic uncertainty (by enhancing parametric knowledge) and improves accuracy (Kaplan et al., 2020), when uncertainty arises, fallback behaviors persist even in the largest, best-trained models. Cases of uncertainty inevitably occur in natural interactions with LLMs, due to knowledge cutoffs, false premises in prompts, or questions about factual information not present in the training data. Ad-hoc fixes, like decoding schemes, fail to address the core issue, often replacing easily detectable degenerate text with subtler and potentially harmful hallucinations. The shift from degenerate text to hallucinations is particularly concerning in the context of scalable oversight (Bowman et al., 2022; Kenton et al., 2024) making model uncertainty harder to detect.
@@
The rest of the paper is organized as follows: Section 2 discusses the different fallback behaviors in LLMs and defines naming conventions, while Section 3 outlines the experimental settings. The empirical results are divided into three parts: Section 4 presents results on the effect of modeling choices, such as parameter count and training procedures, on shifts in fallback behaviors; Section 5 shifts focus to the effect of inference hyperparameters, such as prompting templates and decoding methods, on the emergence of fallbacks; and lastly, Section 6 examines the order of appearance of different fallbacks within a single generation, with a particular focus on text degeneration.
@@ 2 Fallback Behaviors
Since LLMs emerged as powerful generative tools (Radford et al., 2019; Brown et al., 2020), numerous reports have documented their unwanted behaviors. Holtzman et al. (2020) showed that greedy decoding methods can cause degenerate generations, such as incoherent text or repetitions. They propose the use of nucleus sampling to mitigate this problem. In another study, Nasr et al. (2023b) demonstrate that when given out-of-distribution inputs, models might output their training data verbatim. More recently, growing evidence (Zhu et al., 2024; Zhang et al., 2023; Aichberger et al., 2024) suggests that even the best models often generate convincing yet factually incorrect text, typically referred to as “hallucinations” or “confabulations” (Ji et al., 2022; Bang et al., 2023). Xiao & Wang (2021b) investigate model uncertainty, defining it as the probabilistic uncertainty in predicting the next token which is influenced by the model’s training and knowledge, and identify it as a cause of hallucinations. [^3] While the model may experience internal uncertainty when predicting the next token, this uncertainty might not be reflected in the logits. Once a fallback behavior is selected, it may elevate certain tokens, leading to high logit values for these tokens in the final layers.
Figure 2: Larger models resort to more complex fallback behaviors. Pythia Models with larger parameter counts produce more correct facts (green) and hallucinations (orange) while less repeating facts (blue). The green line indicates the number of ground truth answers in Triv.
In this paper, we propose to view these seemingly independent phenomena as fallback behaviors that models exhibit under uncertainty and investigate the relationships between them. Specifically, we focus on epistemic uncertainty, which refers to uncertainty due to a lack of knowledge (Hou et al., 2023). We hypothesize that the strength of an LLM influences its fallback behaviors, and aim to understand what factors determine how a model would behave in cases of uncertainty.
Our study considers the following phenomena: 1. Sequence repetition When models face an inability to produce plausible continuations, they tend to repeat previously generated sequences, which are known to be plausible within the current context. 2. Degenerate text As shown by Holtzman et al. (2020), models can generate degenerate text that is not strictly repetitive but follows a consistent pattern, such as enumerating or repeating sentences with variations in subject entities or attributes. 3. Hallucinations We follow prior work in considering hallucinations as untruthful facts relative to the knowledge the model was exposed to during pretraining and the given context (Liu et al., 2022; Huang et al., 2023; Adlakha et al., 2023; Min et al., 2023a; Zhang et al., 2023). In cases that require the recollection of facts which the model cannot produce, it may fabricate coherent and seemingly plausible yet factually incorrect content.
We additionally looked for cases of verbatim recollection of training samples (Nasr et al., 2023a), but found no evidence of such fallback behavior in our setting (see further discussion in Section 6.1).
@@ 3 Experimental Setting
To directly investigate the relation between fallback behaviors and their contributing factors, our setup exposes models to naturally-occurring controlled cases of epistemic uncertainty. Specifically, we consider tasks that require recalling multiple facts about common topics, without engaging in complex reasoning. The tasks cover both structured outputs in the form of lists and open-ended generation, with queries spanning various difficulty levels, as explained next.
@@
We use the following datasets:
  1. 1.
    TriviaFacts (Triv): We manually curate 95 high-quality domain-diverse questions with a list of answers (list questions) that (a) contains multiple elements, but no more than 15, (b) requires only knowledge that appears very frequently in the web, and (c) includes short elements without multiple correct phrasings. Due to its high-quality and highly controllable setup, we base the bulk of our experiments on this dataset, with minor modifications as relevant to each experiment.
  2. 2.
  3. 3.
    Qampari (Qamp): We sample 100 questions from the dataset introduced by Amouyal et al. (2023), which contains naturally occurring list questions with answers from Wikipedia. [^4] Though similar to our TriviaFacts dataset, this dataset has non-exhaustive answer sets and varying phrasings for the same answers, which can lead to correct answers being flagged as hallucinations.
  4. 4.
    FakeQampari (FQamp): We replace the subjects from Qampari with made-up names, verifying they do not exist. Introducing fictitious subjects reflects real-world queries with false premises (Brahman et al., 2024; Yuan et al., 2024), a common occurrence that forces models into epistemic uncertainty.
Table 1 provides example questions, and additional details on the datasets are in Appendix B. Notably, unlike prior works that study model uncertainty by quantifying it (Xiao & Wang, 2021a; Lin et al., 2023; Zhang et al., 2024), here we introduce model uncertainty by construction. Namely, the above tasks require extensive recall of factual knowledge that is unambiguous and easy to evaluate, such that incorrect answers from the model must be attributed to this uncertainty.
Generating predictions
Notably, in some cases, we prompt the model to generate more facts than actually exist to observe its behavior. While seemingly synthetic, this setup mirrors common scenarios in natural LLM usage, where users prompt models to recall multiple facts, often on topics where the model may have knowledge gaps. These gaps arise from factors like knowledge cutoffs (Kadavath et al., 2022; Hou et al., 2023), false premises in prompts (Brahman et al., 2024), or queries about data not present in the pretraining set, such as proprietary or copyrighted material. We design our testbed to be diverse and representative, while remaining highly controllable, ensuring that epistemic uncertainty is the main cause of any incorrect answers produced by the model. This setup allows us to accurately classify each recalled fact into one of the possible fallback categories.
Evaluation metrics
Given a generated output, we parse it into a set of facts and evaluate each one as correct, repeated, or hallucinated. To avoid overestimating the model’s factuality or hallucination rate, we classify only the first appearance of each fact as correct/hallucination, with all subsequent appearances treated as repetitions. For the list questions datasets (Triv, Qamp and FQamp), this parsing is mostly straightforward as the generations are structured as lists and the ground truth is given as a set of answers. For open-ended generation, we use FactScore Min et al. (2023a) to extract atomic facts and verify them against the entities’ Wikipedia entries. As models frequently continue generating tokens after the completion of the instruction prompt, we further detect what the model did at that point. Namely, we consider the following options: 1) generating EOS token, 2) changing the topic, for example, by creating a new list/biography which we refer to as topic change or 3) continuing to predict tokens indefinitely (until the token budget is exhausted) within the same sentence/paragraph of the answer which we note by bad format. For additional information on the parsing process and example generations and endings, see Appendix C.
Models
We perform our experiments on a variety of model families, sizes, pretraining corpora sizes and finetuning stages. We evaluate both the base and chat-specific checkpoints of Llama 2555We observed that Llama 2 13b, as released on HuggingFace, produces extremely poor results, leading us to suspect an incorrect upload of the weights. To avoid drawing incorrect conclusions based on a potentially defective model, we excluded it from our testbed. Further details and examples are provided in Section E.1. and Llama 3 which were finetuned on instruction and dialogue data  Touvron et al. (2023); Meta AI (2024). Secondly, we use OLMo Groeneveld et al. (2024), which comes in multiple model sizes, includes intermediate checkpoints throughout the pretraining phase and also offers instruction-tuned variants. Finally, we make use of the Pythia model family Biderman et al. (2023) which comes in 8 different scales, each with 154 intermediate pretraining checkpoints. We also include Dolly models which are instruction-tuned Pythia checkpoints (Conover et al., 2023). This suite of models allows us to control for different factors, such as pretraining data, model scale and training procedure, examining their effect on the emergence of fallback behaviors. We used FP16 checkpoints from HuggingFace for all models, except for Llama 2/3 70b which were loaded with FP8 precision.
4 Fallback Behaviors of Different LLMs
Takeaway: Increasing model strength through extra pretraining, more parameters, or instruction-tuning shifts fallback behaviors from simple to complex, i.e., from repetitions to hallucinations.
Scaling up models and training data improves performance and reduces undesired artifacts (Brown et al., 2020). Incorporating instruction-following phases aligns LLMs’ outputs with human preferences (Ouyang et al., 2022). In this section, we investigate how these improvements influence model behavior under uncertainty. We first focus on the trade-off between sequence repetitions and hallucinations, measured as discrete shifts in fallback behaviors. Due to the various forms and broad definition of degenerate text, measuring its appearance is not straightforward. We revisit this in depth in Section 6, demonstrating that the fallback shift is a continuous rather than discrete process.
Table 1: Example questions and answers for our main datasets.
@@ 4.3 Instruction finetuning shifts behaviors
Repeating the experiments from Section 4.1 and comparing models to their instruction-tuned counterparts, we see a similar shift in fallback behavior; instruction-tuned models generate fewer repeated sequences and more hallucinations on average (Figures 21 and 22 in Appendix E depict the results of all instruction tuned models compared to their base versions on TriviaFacts and Qampari, respectively). For the OLMo and Llama 2/3 family which had undergone more finetuning than Dolly checkpoints, the results are much more pronounced, with the hallucinations portion almost doubling between scales. One exception is Llama 3 70b which produced overall less facts, both correct , repetitions and hallucinations, opting to abstain through stopping the generation early or changing the topic. However, among the facts it it did provide, the hallucinations rate remained similar, with a slight increase from 48 While pretrained LLMs are generally unable to generate sequences shorter than what they encountered during training, instruction-tuned models are more likely to produce an EOS token, preempting the generation early and resulting in fewer facts. However, we also note that while instruction finetuning can improve alignment with human preferences, it can sometimes cause finetuned models to diverge more frequently, resulting in some bad format generations.
Figure 3: Larger models hallucinate instead of abstaining. When completing a list of facts about fictitious entities (FQamp), larger Pythia models hallucinate more, while smaller models repeat facts. Models never abstain from giving incorrect facts.
Figure 4: Models that train longer shift to complex fallbacks. As Pythia-6.9B checkpoints see more training tokens (in billions), they produce more hallucinations and fewer repetitions. The green line shows correct answers upper bound in Triv.
@@ 4.4 Similar trends in open-ended generation
To mimic real-world user requests, we use the BioGeneration dataset, sampling completions from a subset of the Pythia model scales with a temperature of $\tau=0.5$. We use FactScore (Min et al., 2023b) to parse each generated biography into atomic facts and let ChatGPT 3.5 (Ouyang et al., 2022) verify them against Wikipedia entries. [^6] While FactScore, which relies on ChatGPT 3.5, can miss some atomic facts or incorrectly label them, we assume such errors occur at similar rates across generations, resulting in reliable trend analysis. Without the limitation on number of facts to produce, we observe that the larger models also generate more facts. Interestingly, when averaging over the very-rare entities (Figure 6) we see that even in this natural scenario, when models are required to elaborate on topics they know little about, they fall back to the same behaviors, with shifts occurring as predictably as in the controlled settings. Similar trends emerge over the rest of the popularity levels, though the more frequent an entity is, the more likely the models are to be able to recall facts for them and less uncertainty they face, thus making the results less useful for our analysis (Figures 35, 35, 37, 37 and 38). This aligns only partially with Min et al. (2023b), who found stronger models generate more atomic facts and struggle more with rare entities. However, they observed that within a model family, larger models are generally more precise (i.e., hallucinate less). In contrast, Lin et al. (2022) found the largest models are the least truthful, which aligns with our findings.